OpenStudy (anonymous):

If you have ever swum in a pool and your eyes The pH level measures the concentration of hydronium ions and can be modeled by the function p(t) = –log10 t. The variable t represents the amount of hydronium ions, and p(t) gives the resulting pH level. Water at 25 degrees Celsius has a pH of 7. Anything that has a pH lower than 7 is called acidic, while a pH above 7 is basic, or alkaline. Seawater has a pH just over 8, while lemonade has a pH of approximately 3. Create a graph of the pH function. Locate on your where the pH value is 0 and where it is 1. You may need to zoom in on your graph.

OpenStudy (anonymous):

I KNOW T= 1/10 , I JUST NEED TO KNOW HOW TO GRAPH IT PLEASE

OpenStudy (aum):

\[ p(t) = -\log_{10}t \] when t = 1, p(t) = 0 when t = 10, p(t) = -1 when t = 100, p(t) = -2

OpenStudy (aum):

when t = 1/10 = 10^(-1), p(t) = +1 when t = 1/100 = 10^(-2), p(t) = +2

OpenStudy (aum):



OpenStudy (anonymous):

Thank you so so much @aum

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

but when i graphed it i put this p(1)=-log(1t) correct? @aum